Circular No. 8460

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004gy Further to IAUC 8454, R. Quimby, C. Gerardy, P. Hoeflich, and J.

C. Wheeler report the discovery of an apparent supernova in unfiltered CCD images taken on 2004 Dec. 28.45 (at mag ∼ 17.0) and 31.38 UT (mag ∼ 17.1). SN 2004gy is located at α = 13^{h}03^{m}58^{s}.42, δ = +26^{o}14′03′′.7 (equinox 2000.0; note that α was given incorrectly on CBET 104), which is 5′′.4 west of the apparent host galaxy. ROTSE-IIIb images taken on Dec. 20.46 (limiting mag ∼ 18.0) showed nothing at this position.

SUPERNOVA 2004gz IN MCG +10-23-45 Further to IAUC 8459, T. Puckett and L. Cox report the discovery

of an apparent supernova (mag 16.4) on an unfiltered CCD frame taken with the 0.35-m automated supernova patrol telescope on 2004 Dec. 28.43 UT; the new object was confirmed at mag 16.5 on frames taken by M. Rice and L. Rice (Mayhill, NM, 0.35-m reflector) on 2005 Jan. 8.47 and at mag 16.3 on a CCD image taken by O. Trondal (Oslo, Norway, 0.25-m reflector) on Jan. 9.2. SN 2004gz is located at α = 16^{h}13^{m}45^{s}.71, δ = +61^{o}01′25′′.9 (equinox 2000.0), which is 1′′.2 east and 2′′.5 north of the center of MCG +10-23-45. The object is not present on images taken by Puckett on 2004 Aug. 19, Sept. 6, or 23 (limiting mag ∼ 20.0).

V1118 ORIONIS P. Williams, Heathcote, N.S.W., reports that this eruptive variable

(cf. IAUC 4676, 4966), usually near mag 18, appears to be undergoing an outburst and is near maximum brightness for the first time since 1997, as indicated by his visual magnitude estimates: 2004 Nov. 14.493 UT, [13.5; Dec. 18.497, [13.5; 2005 Jan. 5.509, 14.0; 5.518, 14.0. CCD images obtained by C. Bembrick (Mt. Tarana, N.S.W.; mag V = 14.0) on Jan. 6.5 and by S. Lee (Coonabarabran, N.S.W.) on Jan. 6.6 confirm the current bright state.

COMET C/2004 Q2 (MACHHOLZ) Total-magnitude and coma-diameter estimates: 2004 Dec. 4.28 UT,

5.3, 16′ (C. S. Morris, Fillmore, CA, 10×50 binoculars); 9.25, 4.7, – (S. J. O’Meara, Volcano, HI, naked eye); 18.33, 4.3, 18′ (Morris; 5^{o}.5 tail in p.a. 28^{o} and 0^{o}.5 fan in p.a. 260^{o}); 25.63, 3.9, 16′ (S. Yoshida, Kanagawa, Japan, 7-mm refractor); 30.86, 3.7, 30′ (T. Karhula, Virsbo, Sweden, naked eye); 2005 Jan. 6.81, 3.5, 22′ (M. Meyer, Kelkheim, Germany, naked eye).

2005 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green